OpenStudy (anonymous):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (anonymous):

f(x) = 3(x + 4)2 + 1

OpenStudy (anonymous):

g(x) = 2x2 −16x + 15

OpenStudy (anonymous):

Can someone help?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

For the first one, you can find the vertex (p,q) because it is in vertex form y=a(x-p)^2+q

OpenStudy (anonymous):

(4,1)

OpenStudy (anonymous):

The second one you have to complete the square: g(x)=2x^2-16x+15 g(x)=2(x^2-8x)+15 g(x)=2[(x^2-8x+16)-16}+15 g(x)=2(x-4)^2-17

OpenStudy (anonymous):

your vertex is wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

well i said the vertex form is y=a(x-p)^2+q, if you plugged in (4,1), the vertex form will be y=3(x-4)^2+1

OpenStudy (anonymous):

So change the sign of the x coordinate

OpenStudy (anonymous):

(-4,1)

OpenStudy (anonymous):

Exactly, did you find the vertex for g(x)?

OpenStudy (anonymous):

the second one is (4,1)7

OpenStudy (anonymous):

(4,17)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

But what about the third one?

OpenStudy (anonymous):

and for the last one, just look fot the coordinate that is either the maximum or minimum

OpenStudy (anonymous):

minimum is (1,-3)

OpenStudy (anonymous):

In this case, the vertex is the minimum value so what is the coordinate that is the lowest point?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the axis of symmetry?

OpenStudy (anonymous):

how do i find that again?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

x=-b/2a

OpenStudy (anonymous):

oh shoot, I re-read your question, i showed you how to find the vertex.

OpenStudy (anonymous):

ok, so let's restart

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you need to convert f(x) into a quadratic equation, so f(x) = 3(x + 4)^2 + 1 f(x)=3(x^2+8x+16)+1 f(x)=3x^2+24x+49

OpenStudy (anonymous):

now that its a quadratic equation (ax^2+bx+c), a is 3 and b is 24

OpenStudy (anonymous):

plug it into the axis of sym. formula -b/2a and you'll have -24/3*2

OpenStudy (anonymous):

that's -24 divided by 6, which is -4, and that is the axis of sym.

OpenStudy (anonymous):

g(x) is already a quadratic equation so what would a and b be?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

err, not exactly, a is 2 and b is -16, so try again

OpenStudy (anonymous):

-4 my bad lol

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

the axis of sym for the graph is pretty easy to find, it's just the x-coordinate of the vertex which we found earlier

OpenStudy (anonymous):

(1,-3)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so wait in order from biggest to smallest would be f(x), g(x) and the graph?

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

well f(x)'s aos = g(x)'s aos so im not sure if you want to specify that they're equal to each other

OpenStudy (anonymous):

but the graph's aos is a positive number while the other two are negative, so do you think it would be the smallest?

OpenStudy (anonymous):

lol, can you show me how to find each axis of symmetry again :)

OpenStudy (anonymous):

dude i spent literally spent 20 minutes explaining, all I can say is x=-b/2a is used to find the aos. As long as the function is quadratic, it is pretty easy to figure it out

OpenStudy (anonymous):

I know and thx so much :)